OpenStudy (anonymous):

x g(x) −2 2 −1 −3 0 2 1 17 Which of the following is a true statement for this function? The function is increasing from x = −2 to x = −1. The function is increasing from x = 0 to x = 1. The function is decreasing from x = −1 to x = 0. The function is decreasing from x = 0 to x = 1.

OpenStudy (anonymous):

I think its A

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

it's the second one, because the output of 0=2 and the output of 1=17, so it is increasing from 0 to 1

OpenStudy (anonymous):

Okay, thanks!

OpenStudy (anonymous):

you're welcome :P

OpenStudy (anonymous):

Water coming out from a fountain is modeled by the function f(x) = −x2 + 6x + 6 where f(x) represents the height, in feet, of the water from the fountain at different times x, in seconds. What does the average rate of change of f(x) from x = 2 to x = 5 represent? The water travels an average distance of 4 feet from 2 seconds to 5 seconds. The water travels an average distance of 1 foot from 2 seconds to 5 seconds. The water falls down with an average speed of 2 feet per second from 2 seconds to 5 seconds. The water falls down with an average speed of 1 foot per second from 2 seconds to 5 seconds. @Nnesha

OpenStudy (anonymous):

make sure to medal XD

OpenStudy (anonymous):

I think the second one is D? I am just trying to check my answers ><

OpenStudy (anonymous):

If D was correct then it would make A true also... i'm pretty sure XD. From 2-5 seconds it would be a total of 4 seconds so if the water it traveling at 1 foot per second for 4 seconds, it would be a total of 4 feet.

OpenStudy (anonymous):

i think

Nnesha (nnesha):

average rate of change means slope the formula to find slope is \[\huge\rm \frac{ y_2 -y_1 }{ x_2 -x_1 }\] given x=2 x=5 you have to find f(5) and f(2) so substitute x for 5 into f(x) function to find y

Nnesha (nnesha):

\[\huge\rm f(x) = −x^2 + 6x + 6\] replace x with 5 and then with 2 solve let me know what you get

OpenStudy (anonymous):

f(x)=61 f(x)=22

Nnesha (nnesha):

nope try again

Nnesha (nnesha):

remember it's x^2 so it would be (2)^2

OpenStudy (anonymous):

So, would the number be negative? I confused now lol

Nnesha (nnesha):

show ur work how did you replaced x with 2 ?

OpenStudy (anonymous):

f(x)=-2^2 +6 x 2 + 6?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

So, then f(x)=-4+6 x 2 + 6?

OpenStudy (anonymous):

I see, they were supposed to remain negative

Nnesha (nnesha):

yes right! so final answer is what

OpenStudy (anonymous):

f(x)=-4 plus 12 plus 6 f(x)=8 plus 6 f(x)=14

Nnesha (nnesha):

yes right now replace x with 5

OpenStudy (anonymous):

f(x)=-25 plus 30 plus 6 f(x) = 5 plus 6 f(x)=11

Nnesha (nnesha):

yes right so order pairs are ( 2,14)(5,11) now plug them into the slope formula

OpenStudy (anonymous):

2^2-5^1 over 14^2 -11^1

Nnesha (nnesha):

ohh wait Y-2 means 2nd y not square

Nnesha (nnesha):

\[\huge\rm (x_1 ,y_1) (1st ~x , 1st ~y)\]

OpenStudy (anonymous):

wait what

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @gigirained 2^2-5^1 over 14^2 -11^1 \(\color{blue}{\text{End of Quote}}\) it's not 14^2

Nnesha (nnesha):